Sana Rosa Company standard rate standard time per unit standard price standard quantity standard rate $24.00 per hr 15 bd. ft Objective 2
media%2F312%2F312500dd-1a36-4dfd-bb72-ca
Answer #1
Std cost of the product:
Material (15 units @ 9.50) 142.5
Labour (3.50 hours @ 24) 84
Variable Overheas (3.50 hours @ 3.20) 11.2
Fixed Overheads (3.50 hours @1.20) 4.2
Std cost of Product 241.9
Q.
Std hours allowed : (2709*2): 540 hours
Actual hours: 580 hours
Std rate per hour: 12
Actual rate per hour: 11.80
Labour rate variance = Actual Hours (Std rate-Actual rate)
580 (12-11.80 ) = 116 F
Labour time variance = Std rate (Std hours-Actual hours)
12 (540-580) = 480 U
Total labour cost variance= Std hours*Std rate -Actual rate*Actual hours
12*540 - 11.80*580 = 364 U
Direct labour cost debited to WIP:
Std hours allowed for actual output 540
Std rate per hour 12
Std cost to be debited to WIP 6480